write a quadratic equation having the given numbers and solutions. -10 and -1. I don't think that this can be done with these given numbers. Help. Answer by ankor@dixie-net.com(22740) 
 Yes it can, think of it as working backwards from solving a quadratic by factoring
:
x=-10, from a factor (x+10)
x=-1, from a factor (x+1)
:
FOIL (x+10)*(x+1) = x^2 + x + 10x + 10
:
y = x^2 + 11x + 10; is the equation with solutions of -10, -1